OpenStudy (anonymous):

a line passes through the point (12, -1) Sum of the intercept made by the axes is 7 what is the equation of the line?

hartnn (hartnn):

can u find the intercepts ?

hartnn (hartnn):

u can refer the question where you had equal intercepts....

OpenStudy (anonymous):

how?

hartnn (hartnn):

for the line y=mx+c to get y-intercept, put x=0 in y=mx+c to get x-intercept, put y=0 in y=mx+c

OpenStudy (anonymous):

after that?

OpenStudy (anonymous):

c + mc=7m ?

hartnn (hartnn):

sum of intercepts = 7 so u get an equation with m and c. use point (12,-1) to find other equation in m and c

hartnn (hartnn):

y- intercept = c \(\checkmark\) x- intercept = ??

OpenStudy (anonymous):

I am not getting it :/

OpenStudy (anonymous):

c/m

hartnn (hartnn):

to get x-intercept, put y=0 in y=mx+c 0=mx+c mx=-c x= -c/m = x-intercept sum = c-c/m = 7 or mc - c = 7m

OpenStudy (anonymous):

ohk and after that?

hartnn (hartnn):

u have point 12,-1 on y=mx+c 12 = -m +c or c= 12+m substitute this c in mc - c = 7m you will get a quadratic in m can u solve quadratic equations ?

OpenStudy (anonymous):

go it :)

OpenStudy (anonymous):

I again got stuck at the quadratic eqn mc-m-19=0 after that?

hartnn (hartnn):

c= 12+m

hartnn (hartnn):

mc - c = 7m c(m-1) = 7m (m+12)(m-1) = 7m m^2 +11m -12 =7m m^2 + 4m -12= 0

OpenStudy (anonymous):

solved it :) Thanks a lot

hartnn (hartnn):

got the values of m and c ?

OpenStudy (anonymous):

m= 2 or -6

hartnn (hartnn):

yup

OpenStudy (anonymous):

which one to take?

hartnn (hartnn):

so u have 2 lines here.....m=2,c=14 m=-6,c=6

hartnn (hartnn):

both the lines satisfy given conditions, so both are correct

OpenStudy (anonymous):

ohk so now?

OpenStudy (anonymous):

ohk

hartnn (hartnn):

y=mx+c y=2x+14 y=-6x+6

hartnn (hartnn):

oh, wait, 12,-1 does not satisfy y=2x+14 2(12)+14 is not -1 -6(12)+6 is still not -1..... i am so sorry!! i made a mistake while substituting the point!

hartnn (hartnn):

u have point 12,-1 on y=mx+c 12 = -m +c <------NO -1=12m+c or c= -1-12m substitute this c in mc - c = 7m you will get a quadratic in m can u solve quadratic equations ?

OpenStudy (anonymous):

i can

hartnn (hartnn):

(-1-12m)(m-1)=7m -12m^2+11m+1=7m 12m^2 - 4m -1=0

OpenStudy (anonymous):

what will be the final equation then?

hartnn (hartnn):

what values of m u got?

OpenStudy (anonymous):

wait I am doing the whole thing again

hartnn (hartnn):

sorry for that mistake.....

OpenStudy (anonymous):

see I am getting the eqn 12m^2 -4m+1=0 is that right?

hartnn (hartnn):

i worked it out to get 12m^2 - 4m -1=0 did u do any sign mistake?

OpenStudy (anonymous):

I did...I corrected it now

OpenStudy (anonymous):

so m= -1/2 or 1/6 now?

OpenStudy (anonymous):

and c= -3 or 5

OpenStudy (anonymous):

is that right?

hartnn (hartnn):

check m=1/2 or -1/6

OpenStudy (anonymous):

check as in? How ? through which eqn?

hartnn (hartnn):

12m^2 - 4m -1=0 m= 1/2 or -1/6 u got it as -1/2 ot 1/6 incorrectly....so i told u to verify again

OpenStudy (anonymous):

see what I did is 12m^2 - 6m + 2m -1=0 is that okay?

hartnn (hartnn):

yes

OpenStudy (anonymous):

(6m+1)(2m-1)?

hartnn (hartnn):

=0 yes

OpenStudy (anonymous):

so its the same what I did before How is it wrong?

hartnn (hartnn):

(6m+1)(2m-1)=0 6m+1=0 or 2m-1=0 m=-1/6 or n=1/2

OpenStudy (anonymous):

yeah now?

hartnn (hartnn):

fin 2 corresponding values of c

OpenStudy (anonymous):

after that?

hartnn (hartnn):

did u get c as 1 or -7 ?

hartnn (hartnn):

m= -1/6, c= 1 ----> y=(-1/6)x+1 m= 1/2, c= -7----->y=x/2-7

OpenStudy (anonymous):

I didn't get the value of c

hartnn (hartnn):

c= -1-12m = -1-12(-1/6) = ?? c= -1-12m =-1-12(1/2) =??

OpenStudy (anonymous):

I got it

hartnn (hartnn):

:)